# Scalable Oversight via Partitioned Human Supervision

### v1

@@ 3 Experiments
In this section, we explain our experiments. We use OpenAI API Platform for all of our experiments.
3.1 Experiment Setup
We evaluate our unbiased and mixed estimators in three settings: (I) Statistical validation on standard multiple-choice benchmarks, where we verify unbiasedness and variance properties (see Eqs. (2), (4), (16)); (II) Practical evaluation on two real-world classification benchmarks: the Japanese financial dataset EDINET-Bench (Sugiura et al., 2025), where each option corresponds to a professional industry domain and annotators provide weak complementary feedback; and the English Medical Abstracts dataset (Schopf et al., 2023), which covers multiple disease categories with natural clinical text. (III) Agentic training in agent-based workflows, where we replace ordinary accuracy with our proposed estimator as the fitness function and examine the resulting agent performance. For agent scaffolding, we use ADAS (Hu et al., 2025) and AFlow (Zhang et al., 2025c). For details of the labeling protocol, please refer to Appendix B.
@@ 3.2 Statistical Validation of Complementary and Mixed Estimators
Statistical validation is conducted on four benchmarks. Since oracle labels are available, complementary labels are synthetically constructed following Appendix B. We report three types of evaluation metrics. (A) Unbiasedness and variance are examined by subsampling 300 ordinary and 300 complementary labels (120 for GPQA due to dataset size). Accuracy is estimated using Eq. (2), and variability is quantified through standard deviations, reported as $\widehat{A}\pm\widehat{\sigma}$ with $\widehat{\sigma}$ denoting the estimated standard deviation. (B) The relative efficiency of complementary labels is assessed by computing the number of complementary samples required to achieve the same variance as $n_{\mathrm{o}}$ ordinary labels, according to Eq. (4), where the full ordinary dataset is used to approximate the ground-truth accuracy $A$. (C) To mimic practical annotation protocols (Appendix B), we assume that each ordinary label can be replaced by (K1)(K-1) complementary labels, where $K$ is the number of answer options. This setting is used to evaluate the mixed estimators in Eq. (7) and Eq. (9).
Each dataset is evaluated on all five metrics using gpt-5-nano (OpenAI, 2025b), with three independent runs averaged; aggregated outcomes are summarized in Table 1. For the MATH-MC benchmark, we use GPT-4.1-nano instead of GPT-5-nano. This choice is due to (i) output-format constraints in GPT-5-nano that conflict with our final-answer extraction protocol, and (ii) the need to avoid ceiling effects: in our pilot runs, GPT-5-nano attains near-ceiling accuracy on MATH-MC, which compresses the observable gaps among estimators.
Table 1: Performance of estimators across benchmarks. Values are reported as mean accuracy $\pm$ standard deviation across three random seeds; the average of per-run standard deviations is additionally shown in parentheses. Detailed estimator setups and confidence-interval analyses are provided in Appendix F. The last column reports the macro-average across datasets (counting Math and Math-COT separately). For each benchmark, the estimator whose mean accuracy is closest to the Ord-Eval reference is highlighted in bold.
@@
Among mixture estimators, IVW consistently improves over the equal-weighted baseline in within-run standard deviation (IVW-0.5), confirming that inverse-variance weighting provides the optimal linear combination of ordinary and complementary labels. ML further achieves narrower intervals owing to its joint likelihood formulation. However, ML requires more precise pilot estimates of accuracy and may thus be less robust in practice, whereas IVW remains stable by directly incorporating variance estimates. This distinction is supported by the within-run standard deviations, which indicate that IVW achieves competitive efficiency with simpler computation.
@@
(a) Optimal $w$ as a function of the number of samples (normalized by 300).
(b) Corresponding variance to the given $w$ on a log scale.
To further examine whether the closed-form coefficient yields the optimal trade-off, we conduct an ablation on MedQA-USMLE by varying the number of complementary labels from $n_{\mathrm{c}}=n_{\mathrm{o}}$ to $n_{\mathrm{c}}=20n_{\mathrm{o}}$ (with $n_{\mathrm{o}}=300$). Results are averaged over three random seeds. As shown in Figure 2, when $n_{\mathrm{c}}\approx 3n_{\mathrm{o}}$—the setting used in Table 1—the optimal weight is close to the fixed baseline $w=0.5$. However, for $n_{\mathrm{c}}<n_{\mathrm{o}}$ or $n_{\mathrm{c}}>6n_{\mathrm{o}}$, the optimal variance is consistently lower than the baseline. This behavior is intuitive: once the number of complementary labels exceeds the variance-matching threshold in Eq. (4), the complementary estimator alone has smaller variance, and the closed-form IVW solution naturally shifts weight away from ordinary labels. By contrast, the fixed baseline cannot adapt, leading to consistently higher variance.
@@ 3.3 Benchmark Evaluation in Real-World Tasks
We evaluate on two benchmarks. First, we use EDINET-Bench (Sugiura et al., 2025), a Japanese financial report classification dataset, where each industry corresponds to a professional domain, and we adopt the prompting setup of Sugiura et al. (2025). Since the original dataset contains only 496 samples (comparable to GPQA), we extend it following their released code, resulting in 3,269 samples; we refer to this as EDINET-Bench Extended. To our knowledge, there are no public medical benchmarks in which each class explicitly corresponds to a medical specialization. We therefore adopt the Medical Abstracts dataset (Schopf et al., 2023), which consists of paper abstracts labeled by disease categories, simulating the task of narrow domain experts classifying diseases.
@@
Table 2: Mean acc. $\pm$ standard deviation across 3 seeds, w/ the average of per-run standard deviations. Bold: Lowest variance across 3 runs. See Appendix. F for details.
@@ 3.4 Agent Search with Estimator Feedback
Finally, we show that weak human supervision can be leveraged not only for evaluation but also as training signal for AI systems. We evaluate agentic flows under complementary-label signals on three benchmarks. GPQA serves as the most challenging setting in our study (Table 1); Math-MC focuses on reasoning ability; and Medical Abstracts best captures our practical scenario. Following Hu et al. (2025), we include CoT (Wei et al., 2022), CoT-SC (Wang et al., 2023), Self-Refine (Madaan et al., 2023), LLM Debate (Du et al., 2024), Step-back Abstraction (Zheng et al., 2024), Quality-Diversity (Lu et al., 2025), and Role Assignment (Xu et al., 2025) as manually designed baselines, and compare against ADAS (Hu et al., 2025) and AFlow (Zhang et al., 2025c). We define the accuracy computed by $q$ as complementary accuracy and apply Eq. (2) to obtain the transformed complementary accuracy. For ADAS and AFlow, we follow the original protocol (see Appendix G), computing both complementary and transformed complementary accuracy on the validation set and reporting the better-performing variant.
@@
Figure 3 compares ADAS and AFlow against strong baselines on GPQA, Math-MC, and Medical Abstracts. Both agentic systems consistently outperform manually designed workflows once weak complementary-label signals are incorporated, demonstrating that weak human feedback can provide learning signal. While the choice of estimator (raw vs. transformed complementary accuracy) leads to different outcomes depending on task difficulty, the overall pattern remains robust: weak signals suffice to improve performance beyond established baselines, with additional variance analysis provided in Appendix H. These findings reinforce the central theme of our study: humans’ weak labels can provide a reliable and scalable supervisory channel for steering agentic AI systems.
@@ 4 Related Work
LLMs have inspired a wide range of system architectures (Wei et al., 2022; Wang et al., 2023; Madaan et al., 2023; Shinn et al., 2023b; a), in particular by incorporating multiple LLMs into collaborative frameworks (Du et al., 2024; Liu et al., 2024; Chen et al., 2024; Wu et al., 2024; Wang et al., 2024c). Early studies relied on manually designed agentic workflows, which demanded substantial manual design and domain expertise. More recent work has shifted toward automated agent discovery, in which LLMs themselves are used to generate, refine, and optimize workflow structures (Hu et al., 2025; Zhang et al., 2025a; b; c), but required ground-truth supervision to construct a fitness score. We investigate whether automated agentic system design can proceed even with weaker training signals, i.e., complementary labels.
@@ 5 Conclusion
Appendix A Connection between the complementary-label estimator and complementary loss
@@
Appendix B Data Labeling and Estimation Protocol
In multiple-choice evaluation, the effective “class” is the index position of an option rather than its textual content. To satisfy the uniformity condition in Eq. (1), the labeling process proceeds as follows. For each item, all options are first shuffled uniformly at random. After shuffling, the correct index $y$ is recorded. A complementary label is then created by sampling a wrong index $\bar{y}$ uniformly from the K1K-1 remaining positions. The final published datum is $(x,\text{shuffled options},\bar{y})$. When the ground truth is known, this procedure can be automated: $y$ is only used internally for generating $\bar{y}$, and the released dataset never exposes $y$. When labels are collected from human raters, the annotation interface can enforce the same randomization in the backend to guarantee uniformity. Any deviation, such as asking raters to pick the ‘most plausible wrong option,’ would violate the assumption behind Eq. (1). Exploring how to deal with alternative data or label collection mechanisms is left for future work.
@@ Appendix B Data Labeling and Estimation Protocol
11:Compute $S_{\mathrm{o}}=\sum Z$, $S_{\mathrm{c}}=\sum W$.
@@
Appendix C One-step Newton Update and IVW Equivalence
@@
Appendix D Proof of Theorem 4
@@
Appendix E Practical plug-in forms of the Bernstein bound
@@
Appendix F Estimator Configurations

### v2

@@ 3 Experiments
We evaluate our unbiased and mixed estimators in three settings: (I) Statistical validation on standard multiple-choice benchmarks, where we verify unbiasedness and variance properties (see Eqs. (2), (4), (16)); (II) Practical evaluation on two real-world classification benchmarks: the Japanese financial dataset EDINET-Bench (Sugiura et al., 2026), where each option corresponds to a professional industry domain and annotators provide weak complementary feedback; and the English Medical Abstracts dataset (Schopf et al., 2023), which covers multiple disease categories with natural clinical text. (III) Agentic training in agent-based workflows, where we replace ordinary accuracy with our proposed estimator as the fitness function and examine the resulting agent performance. For agent scaffolding, we use ADAS (Hu et al., 2025) and AFlow (Zhang et al., 2025b). For details of the labeling protocol, please refer to Appendix C.
@@ 3.1 Statistical Validation of Complementary and Mixed Estimators
Statistical validation is conducted on four benchmarks. Since oracle labels are available, complementary labels are synthetically constructed following Appendix C. We report three types of evaluation metrics. (A) Unbiasedness and variance are examined by subsampling 300 ordinary and 300 complementary labels (120 for GPQA due to dataset size). Accuracy is estimated using Eq. (2), and variability is quantified through standard deviations, reported as $\widehat{A}\pm\widehat{\sigma}$ with $\widehat{\sigma}$ denoting the estimated standard deviation. (B) The relative efficiency of complementary labels is assessed by computing the number of complementary samples required to achieve the same variance as $n_{\mathrm{o}}$ ordinary labels, according to Eq. (4), where the full ordinary dataset is used to approximate the ground-truth accuracy $A$. (C) To mimic practical annotation protocols (Appendix C), we assume that each ordinary label can be replaced by (K1)(K-1) complementary labels, where $K$ is the number of answer options. This setting is used to evaluate the mixed estimators in Eq. (7) and Eq. (9).
Each dataset is evaluated on all five metrics using gpt-5-nano (OpenAI, 2025a), with three independent runs averaged; aggregated outcomes are summarized in Table 1. For the MATH-MC benchmark, we use GPT-4.1-nano instead of GPT-5-nano. This choice is due to (i) output-format constraints in GPT-5-nano that conflict with our final-answer extraction protocol, and (ii) the need to avoid ceiling effects: in our pilot runs, GPT-5-nano attains near-ceiling accuracy on MATH-MC, which compresses the observable gaps among estimators.
Table 1: Performance of estimators across benchmarks. Values are reported as mean accuracy $\pm$ standard deviation across three random seeds; the average of per-run standard deviations is additionally shown in parentheses. A brief description of the estimators is provided in the legend below Detailed estimator setups and confidence-interval analyses are provided in Appendix I. The last column reports the macro-average across datasets (counting Math and Math-COT separately). For each benchmark, the estimator whose mean accuracy is closest to the Ord-Eval reference is highlighted in bold. Column-wise deviations from the oracle ($\Delta$) are reported in Appendix L.
@@
Legend:
Ord = ordinary-label estimator; Comp-$n_{o}$ = complementary labels with the same sample size as ordinary; Comp-Var = complementary labels with variance-matched sample size; IVW-0.5 = fixed 0.5/0.5 weighting of ordinary and complementary estimates; IVW = inverse-variance–weighted combination with plug-in weights; ML = joint maximum-likelihood estimator using both label types; Ord-Eval = oracle accuracy on the full dataset.
@@
(a) Optimal $w$ vs. number of samples.
(b) Corresponding variance on a log scale.
Mixture estimators (IVW and ML) reduce the gap to the Ord-Eval reference compared with both ordinary-only and complementary-only estimators across most benchmarks. Among mixture estimators, IVW consistently improves over the equal-weighted baseline in within-run standard deviation (IVW-0.5), confirming that inverse-variance weighting provides the optimal linear combination of ordinary and complementary labels. ML further achieves narrower intervals owing to its joint likelihood formulation. However, ML requires more precise pilot estimates of accuracy and may thus be less robust in practice, whereas IVW remains stable by directly incorporating variance estimates. This distinction is supported by the within-run standard deviations, which indicate that IVW achieves competitive efficiency with simpler computation.
@@
To further examine whether the closed-form coefficient yields the optimal trade-off, we conduct an ablation on MedQA-USMLE by varying the number of complementary labels from $n_{\mathrm{c}}=n_{\mathrm{o}}$ to $n_{\mathrm{c}}=20n_{\mathrm{o}}$ (with $n_{\mathrm{o}}=300$). Results are averaged over three random seeds. As shown in Figure 2, when $n_{\mathrm{c}}\approx 3n_{\mathrm{o}}$ (the setting used in Table 1), the optimal weight is close to the fixed baseline $w=0.5$. However, for $n_{\mathrm{c}}<n_{\mathrm{o}}$ or $n_{\mathrm{c}}>6n_{\mathrm{o}}$, the optimal variance is consistently lower than the baseline. This behavior is intuitive: once the number of complementary labels exceeds the variance-matching threshold in Eq. (4), the complementary estimator alone has smaller variance, and the closed-form IVW solution naturally shifts weight away from ordinary labels. By contrast, the fixed baseline cannot adapt, leading to consistently higher variance.
@@ 3.2 Benchmark Evaluation in Real-World Tasks
We evaluate on two benchmarks. First, we use EDINET-Bench (Sugiura et al., 2026), a Japanese financial report classification dataset, where each industry corresponds to a professional domain, and we adopt the prompting setup of Sugiura et al. (2026). Since the original dataset contains only 496 samples (comparable to GPQA), we extend it following their released code, resulting in 3,269 samples; we refer to this as EDINET-Bench Extended. To our knowledge, there are no public medical benchmarks in which each class explicitly corresponds to a medical specialization. We therefore adopt the Medical Abstracts dataset (Schopf et al., 2023), which consists of paper abstracts labeled by disease categories, simulating the task of narrow domain experts classifying diseases.
@@
Table 2: Mean acc. $\pm$ standard deviation across 3 seeds, w/ the average of per-run standard deviations. Bold: Lowest variance across 3 runs. See App. I for details.
@@ 3.3 Agent Search with Estimator Feedback
Finally, we show that weak human supervision can be leveraged not only for evaluation but also as training signal for AI systems. We evaluate agentic methods under complementary-label signals on three benchmarks. GPQA serves as the most challenging setting in our study (Table 1); Math-MC focuses on reasoning ability; and Medical Abstracts best captures our practical scenario. Following Hu et al. (2025), we include CoT (Wei et al., 2022), CoT-SC (Wang et al., 2023), Self-Refine (Madaan et al., 2023), LLM Debate (Du et al., 2024), Step-back Abstraction (Zheng et al., 2024), Quality-Diversity (Lu et al., 2025), and Role Assignment (Xu et al., 2025) as manually designed baselines, and compare against ADAS (Hu et al., 2025) and AFlow (Zhang et al., 2025b). We define the accuracy computed by $q$ as complementary accuracy and apply Eq. (2) to obtain the transformed complementary accuracy. For ADAS and AFlow, we follow the original protocol (see Appendix J), computing both complementary and transformed complementary accuracy on the validation set and reporting the better-performing variant.
@@
Figure 3 compares ADAS and AFlow against strong baselines on GPQA, Math-MC, and Medical Abstracts. Both agentic systems consistently outperform manually designed workflows once weak complementary-label signals are incorporated, demonstrating that weak human feedback can provide learning signal. While the choice of estimator (raw vs. transformed complementary accuracy) leads to different outcomes depending on task difficulty, the overall pattern remains robust: weak signals suffice to improve performance beyond established baselines, with additional variance analysis provided in Appendix K. To further illustrate the resulting agentic system’s behavior, we present a case study in Appendix M. These findings reinforce the central theme of our study: humans’ weak labels can provide a reliable and scalable supervisory channel for steering agentic AI systems.
@@ 4 Related Work
LLMs have inspired a wide range of system architectures (Wei et al., 2022; Wang et al., 2023; Madaan et al., 2023; Shinn et al., 2023), in particular by incorporating multiple LLMs into collaborative frameworks (Du et al., 2024; Liu et al., 2024; Chen et al., 2024; Wu et al., 2024; Wang et al., 2024c). Early studies relied on manually designed agentic workflows, which demanded substantial manual design and domain expertise. More recent work has shifted toward automated agent discovery, in which LLMs themselves are used to generate, refine, and optimize workflow structures (Hu et al., 2025; Zhang et al., 2025a; 2026; b), but required ground-truth supervision to construct a fitness score. We investigate whether automated agentic system design can proceed even with weaker training signals, i.e., complementary labels.
5 Discussion & Limitations
Assumptions and extensions
The assumption of uniformly querying an index plays a key role in obtaining unbiased complementary labels, and in our labeling protocol this assumption is guaranteed for the practical scenario we consider. However, to make the whole system more robust, we should also consider possible issues that may arise in practice, such as biased answers from experts, uncertain answers (which can be reframed as a type of noise), and overlaps between expert fields. For these cases, a wide range of complementary-label variants has been studied, such as biased complementary labels (Yu et al., 2018; Wang et al., 2024a), noisy complementary labels (Ishiguro et al., 2022), and multiple complementary labels (Feng et al., 2020). We believe that incorporating these formulations would further enhance the robustness of our framework.
Beyond the assumptions on the labels themselves, we also need to consider the structural assumption that the task is posed in a multiple-choice format. The multiple-choice format still has wide application, and in scenarios beyond the standard multiple choice setting, some open-ended questions can be transformed into structured multiple-choice forms to allow the use of our framework. For example, in debugging, modern systems strongly rely on cooperation among different sectors. The end-to-end service may look open-ended, but it can often be separated into different sectors or a hierarchical structure. In particular, when the root cause is unknown, we can ask engineers responsible for different parts of the product to provide weak signals, for example, confirming that the bug does not originate from their component. This kind of knowledge fragmentation provides opportunities to apply our framework.
Limitations
Although our framework has shown its potential for addressing the expertise problem in evaluating foundation models, we would still like to discuss its current limitations. First, the framework assumes the availability of complementary labels. When even complementary labels are scarce or inapplicable, for example, in open-ended problems that are hard to decompose into a multiple-choice pipeline, complementary labels may be ill-defined, and the formulation may need to be revised. Second, when feedback from humans is no longer available and the boundary of the ground truth becomes ambiguous, neither complementary labels nor ordinary labels can be reliably defined. In such cases, new measures of accuracy would be required to revise the framework. We view these current limitations as interesting directions for future research.
@@ 6 Conclusion
Acknowledgements
RY was supported by the RIKEN-AIP Undergraduate Research Program, TI was supported by JST ASPIRE Grant Number JPMJAP25B1, and MS was supported by JST ASPIRE Grant Number JPMJAP2405.
Ethics statement
Our work is motivated by AI safety: we aim to improve oversight of advanced AI systems by enabling evaluation and training even on tasks beyond any single human expert’s competence. This approach is intended as a step toward improving the reliability and alignment of future AI models with human interests, which we believe is a positive contribution to society.
Currently, we do not foresee immediate direct negative social impacts from the research itself. However, we recognize that alignment and oversight techniques can be a double-edged sword: they are dual-use in the sense that they could be repurposed by malicious actors to optimize systems toward harmful objectives (Machine Learning Street Talk, 2025), and even when used defensively, they can be vulnerable to strategic behavior that produces a false sense of safety (Hubinger et al., 2024; Greenblatt et al., 2024). These risks are not unique to our approach, but we caution against such misuse and vulnerabilities.
Reproducibility Statement
Code and scripts to reproduce our experiments are available in our public GitHub repository: Towards Scalable Oversight via Partitioned Human Supervision.
For ADAS (Hu et al., 2025) and AFlow (Zhang et al., 2025b), we follow their official implementations; see Appendix J for additional details. Because our experiments use the OpenAI API, exact replication may vary over time if models are updated or retired. We therefore provide full implementation details, dataset references, and evaluation scripts to support reproducibility as closely as possible.
Appendix A Proof of Corollary 1
Proof.
For a single item, under Eq. (1),
$=\Pr(\hat{Y}=Y)\cdot 1+\Pr(\hat{Y}\neq Y)\cdot\Big(1-\tfrac{1}{K-1}\Big)=\frac{A+K-2}{K-1}.$
If we consider $\hat{q}=\tfrac{1}{n_{\mathrm{c}}}\sum_{i=1}^{n_{\mathrm{c}}}w_{i}$, an empirical version of $\mathbb{E}[W]$, this $\hat{q}$ is an unbiased estimator of $\mathbb{E}[W]$. Hence, $\widehat{A}_{\mathrm{comp}}=(K-1)\hat{q}-(K-2)$ is an unbiased estimator of $A$. ∎
Appendix B Connection between the complementary-label estimator and complementary loss
@@
Appendix C Data Labeling and Estimation Protocol
In multiple-choice evaluation, the effective “class” is the index position of an option rather than its textual content. To satisfy the uniformity condition in Eq. (1), the labeling process proceeds as follows. For each item, all options are first shuffled uniformly at random. After shuffling, the correct index $y$ is recorded. A complementary label is then created by sampling a wrong index $\bar{y}$ uniformly from the K1K-1 remaining positions. The final published datum is $(x,\text{shuffled options},\bar{y})$. When the ground truth is known, this procedure can be automated: $y$ is only used internally for generating $\bar{y}$, and the released dataset never exposes $y$. When labels are collected from human raters, the annotation interface can enforce the same randomization in the backend to guarantee uniformity. Any deviation, such as asking raters to pick the “most plausible wrong option,” would violate the assumption behind Eq 1.
@@ Appendix C Data Labeling and Estimation Protocol
11:Compute $S_{\mathrm{o}}=\sum_{j}z_{j}$, $S_{\mathrm{c}}=\sum_{i}w_{i}$.
@@
Appendix D One-step Newton Update and IVW Equivalence
Setup and notation.
@@ Appendix D One-step Newton Update and IVW Equivalence
Appendix E Proof of Theorem 2
Proof.
Hoeffding part. For i.i.d. $w_{i}\!\in\![0,1]$, Hoeffding gives
[…]
Empirical Bernstein part. Apply the two-sided empirical Bernstein inequality (Maurer and Pontil, 2009) to $\hat{q}=\tfrac{1}{n_{\mathrm{c}}}\sum_{i=1}^{n_{\mathrm{c}}}w_{i}$ with $w_{i}\in[0,1]$:
[…]
Appendix F Proof of Theorem 4
@@
Appendix G Practical plug-in forms of the Bernstein bound
@@
Formally, Eq. (10) is valid for any fixed choice of the weight $w\in[0,1]$ (e.g., a constant choice or an estimate obtained from an independent pilot split). If instead $w$ is chosen adaptively from the same evaluation data (e.g., the plug-in $\widehat{w}_{\mathrm{IVW}}$), the fixed-weight assumption is violated and the bound is no longer guaranteed. In such cases one can (i) revert to the union-bound inequalities of the main text, which hold for arbitrary (possibly data-dependent) $w$ via a triangle inequality; (ii) restore the fixed-$w$ condition through sample splitting or cross-fitting (Chernozhukov et al., 2018; Wager and Athey, 2018); or (iii) restrict $w$ to a finite grid $G\subset[0,1]$ and apply a union bound with $\delta$ adjusted by $|G|$ (Boucheron et al., 2013; Shalev-Shwartz and Ben-David, 2014).
Appendix H Benchmark Details
Appendix I Estimator Configurations